OpenStudy (anonymous):

Trying to solve -6+x/4=-5. Please show me the steps.

OpenStudy (anonymous):

remove the fraction first by multiplying each term by 4.

OpenStudy (anonymous):

is it \(6+\frac{x}{4}=-5\) or \(\frac{6+x}{4}=-5\) ?

OpenStudy (anonymous):

@kperk71 still there?

OpenStudy (anonymous):

I've done that and still don't get the answer

OpenStudy (anonymous):

can you show me your solution? i'll check

OpenStudy (anonymous):

it is -6+x/4=-5

OpenStudy (anonymous):

x=-14

OpenStudy (anonymous):